Acme Oscillators is considering an investment project that has the following rather unusual cash flow pattern.

Year

Cash Flow

0

​$100

1

−460

2

791

3

−602.6

4

171.6

a. Calculate the​ project's NPV at each of the following discount​ rates:

0​%,

5​%,

10​%,

20​%,

30%​,

40​%,

50​%.

b. What do the calculations tell you about this​ project's IRR? The IRR rule tells managers to invest if a​ project's IRR is greater than the cost of capital. If Acme​ Oscillators' cost of capital is

8​%,

should the company accept or reject this​ investment?

c. Notice that this​ project's greatest NPVs come at very high discount rates. Can you provide an intuitive explanation for that​ pattern?
Answer #1

As per rules I am answering the first 4 subparts of the question

1: NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= 100-460/(1+0%)^1+791/(1+0%)^2-602.6/(1+0%)^3+171.6/(1+0%)^4

= 0

2: NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= 100-460/(1+5%)^1+791/(1+5%)^2-602.6/(1+5%)^3+171.6/(1+5%)^4

= $-0.0077

3: NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= 100-460/(1+10%)^1+791/(1+10%)^2-602.6/(1+10%)^3+171.6/(1+10%)^4

=0

4: NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= 100-460/(1+20%)^1+791/(1+20%)^2-602.6/(1+20%)^3+171.6/(1+20%)^4

= 0